27) Consider the following information Quantity Consumed Candy Bars Apples Total Marginal per WeekityUtility Total Utility Marginal Utility 0 10 18 25 29 31 30 0 15 24 30 32 31 29 a. Complete the columns containing the marginal utilities of each good. b. With unlimited income, a consumer described by the above information would consume candy bars andapples. If candy bars cost 25 cents each and apples are two for a quarter (12.5 cents each) and the consumer has $1.50 to spend on candy and apples, what quantity of candy bars and apples will allow the above individual to reach consumer optimum? c.
Answer #1

a)

                       Candy                                                                                            Apples                             

Q

TU

MU

TU

MU

0

0

0

0

1

10

-

15

-

2

18

8

24

9

3

25

7

30

6

4

29

4

32

2

5

31

2

31

-1

6

30

-1

29

-2

b)

Would consume 5 Candies and 4 apples.

c)

MUc/Pc = MUa/Pa

4/25 = 2/12.5 = 0.16

Hence reaching equilibrium, 4 units of each commodity shall be consumed.